Ray and Joan can't decide between two kinds of tiles for their rectangular kitchen floor; either cork tiles that measure 10cmx10cm and cost $0.29 each or ceramic tiles that measure 20cmx50cm costing $2.49 each. The floor measures 4m by 2m. How many dollars do they save by choosing the cheaper tile?
Answer by richwmiller(17219)

20*50=1000 cm^2
10*10=100 cm^2
2m=200 cm
4m=400 cm
200*400 =80000 cm^2
ceramic 80000/1000=80*2.49
cork 80000/100=800*.29
subtract the two